Determine two consecutive multiples of 3 whose product is 270.
Answer by mananth(16949)

Let the numbers be 3x & 3x+3
3x(3x+3)=270
/3
x(3x+3)=90
3x^2+3x=90
/3
x^2+x-30=0
x^2+6x-5x-30=0
(x+6)(x-5)=0
x=-6
x=5
plug the value of x in 3x & 3x+3
the numbers are 15 & 18
OR -15,-18